Reports 1-1 of 1 Clear search Modify search
AdV-TCS (CO2 laser projector)
lorenzini, nardecchia, rocchi, gherardini - 0:20 Tuesday 07 March 2023 (59155) Print this report
NI CO2 actuators alignement check

Today shift has been devoted to the final checks of the NI CO2 actuators on the CP, performing an absorption measurement with gentle unlock.

After 1 hr in CARM Null, at 16.58 UTC, HWS-DET acquisition started (20230306T1757), at 17.02 UTC the ITF unlocked in gentle manner (WF 9).

The PA appeared immediately in the HWS map (fig.1) and the coordinates of the YAG have been evaluated considering the WFs from 100 to 173 (see table below). An example of the YAG OPL in the HWS-DET map, after removing the PA, is shown in fig.2.

NI

 

X [pixel]

Y [pixel]

YAG (230306T1757) 

WFs (100-173)

BARY

426+/-21

518+/-14

MAX

426 +/-20

519+/-14

Then, we checked the CO2 actuator positions, but the CH OPL appeared stretched  in the vertical direction.
Despite the elongated shape, a small horizontal misalignment has been observed, therefore, at 21.00 UTC, the NI HM has been moved in TY BACKWARD by 1200 STEPS (to be conservative)
.

The final CH coordinates are summarized in the table below:

NI

X [pixel]

Y [pixel]

CH 

(230306T2206) 

WFs (30-70)

470 +/-26

522 +/- 23

At the end, the CO2 nominal powers have been restored.

Images attached to this report
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×